Locating the receiver for a radio telescope.
The U.S. Naval research Laboratory designed a giant radio telescope weighting 3450 tons. Its parabolic dish has a diameter of 300 feet and a depth of 44 feet
(a) find an equation in the form y=ax^2 that describes a cross section of this dish.
(b) if the receiver is located at the focus, how far should it be from the dish(vertex)?

Answer by stanbon(75887)

The U.S. Naval research Laboratory designed a giant radio telescope weighting 3450 tons. Its parabolic dish has a diameter of 300 feet and a depth of 44 feet
--------------
(a) find an equation in the form y=ax^2 that describes a cross section of this dish.
Draw the parabolic cross section on a coordinate system.
Vertex at (0,-44); x-intercepts at (-150,0) and (150,0)
Form: (x-h)^2=4p(y-k)
h=0; k=-44; x=150; y=0
150^2=4p(44)
p=127.84
EQUATIOn:
x^2=4(127.84)(y+44)
y=(1/511.36)x^2 - 44
y=0.0019555695...x^2 -44
----------------------------
(b) if the receiver is located at the focus, how far should it be from the dish(vertex)?
Focus at (0,p) = (0,127.84 ft.)
127.84 ft above the vertex
-------------
Cheers,
Stan H.